How many miles of O2 are produced from 5.0 miles of V2O5?
1
Expert's answer
2020-06-22T09:15:18-0400

V2O5 thermal decomposition reaction:

2 V2O5 → 4VO2 + O2

n(O2)=n(V2O5)/2=5.0/2=2.5 (moles)